# CARTE: pretraining and transfer for tabular learning

### v1

@@ Abstract
Pretrained deep-learning models are the go-to solution for images or text. However, for tabular data the standard is still to train tree-based models. Pre-training or transfer is a huge challenge as in general tables have columns about different quantities and naming conventions that vary vastly across sources. Data integration tackles correspondences across multiple sources: schema matching for columns, and entity matching for entries. We propose a neural architecture that does not need such matches. As a result, we can pretrain it on background data that has not been matched. The architecture –CARTE for Context Aware Representation of Table Entries– uses a graph representation of tabular (or relational) data to process tables with different columns, string embeddings of entries and columns names to model an open vocabulary, and a graph-attentional network to contextualize entries with column names and neighboring entries. An extensive benchmark shows that CARTE facilitates learning, outperforming a solid set of baselines including the best tree-based models. CARTE also enables joint learning across tables with unmatched columns, enhancing a small table with bigger ones. CARTE opens the door to large pretrained models embarking information for tabular data.
@@ 1 Introduction
The wide availability of pre-trained models has greatly facilitated machine learning for various data modalities, for example with images (Simonyan & Zisserman, 2015) or texts (Devlin et al., 2019). These models can be downloaded from model hubs, embarking a lot of implicit information and transformations that unleashes the power of deep learning even on small datasets. This paradigm has led to the revolution of foundation models (Bommasani et al., 2021) such as large language models (Touvron et al., 2023). But this revolution has not happened for tabular data despite its huge importance for enterprise and institutional data. One roadblock is that integrating data from tables in the wild is often difficult, sometimes impossible. Different tables might not have any related data and when they do, data integration is a whole field of database research (Doan et al., 2012). One might need to solve correspondence across columns –schema matching– or across data sources with different naming conventions for entries –entity matching.
For lack of matching schemas and entities, pretraining across tables in the wild has so far not been possible. Lack of pretraining reduces the benefit of deep architectures, and applications often rely on tree-based methods (Grinsztajn et al., 2022). Here we introduce a learning architecture that enables bypassing schema and string matching by using a graph representation of tables as well as embeddings for all symbols (column names, table entries). The architecture, dubbed CARTE (Context-Aware Representation of Table Entries), is pretrained on a large knowledge base, to capture information on a vast amount of entities and relations. It can then be fine-tuned on a given downstream task, helping learning even in few shot settings. It can also be used for joint learning across multiple tables, enriching a target table with weakly related sources.
In section 2 we present related work; in section 3 we describe the CARTE architecture and training procedures; and in section 4 we ran an extensive empirical study across many tabular datasets, benchmarking the settings of a single downstream table as well as multiple related ones.
@@ 2 Related Works
Tables are central to many applications. As a result, numerous deep learning methods tailored for this modality have been proposed (Abutbul et al., 2020; Arik & Pfister, 2020; Popov et al., 2019; Gorishniy et al., 2023b; Somepalli et al., 2021). However, their superiority over tree-based methods has been challenged (Grinsztajn et al., 2022; Shwartz-Ziv & Armon, 2021) especially when accounting for speed or robustness (Gardner et al., 2022). While this debate is ongoing, with McElfresh et al. (2023) claiming that neural networks perform well on certain types of tables, and promising architectures recently published (Gorishniy et al., 2023a; Chen et al., 2023), the difficulty of making significant improvements over tree-based methods suggests that deep learning must bring something more to the fight, such as background knowledge.
@@
Most transfer learning has studied the “conventional” settings: transferring across datasets with the same features i.e. columns. Somepalli et al. (2021) demonstrates pre-training on a larger unlabeled version of the table to improve prediction, while Levin et al. (2023) argues that transfer learning bridges the gap between deep learning and tree-based models in settings where there are few data points but large related datasets, such as in their medical setting. In addition, they consider new or missing features in the downstream table, though this requires exact matching between most features.
@@
TabPFN (Hollmann et al., 2023) has made an interesting step in the direction of pre-trained models for tabular learning: it uses a transformer model pre-trained on large amounts of synthetic data to capture the inductive biases of tabular data, leading to very strong performance on small datasets, though it has no dedicated handling for categorical columns, a challenge of tables where trees historically shine.
@@
One challenge of tables –much more tackled by the database literature than the machine learning one– is that many of the entries are discrete, represented as strings. Cerda & Varoquaux (2022) created string-based representations that facilitate learning. The TableVectorizer in Skrub (2024) uses these heuristically to turn columns of different types into numerical matrices well suited for learning. Cvetkov-Iliev et al. (2023) contributed KEN, another approach to embed table entities, closer to our goals of pretraining. They create embeddings of all entities in a knowledge graph, capturing the information in a source such as Wikipedia. These embeddings facilitate learning, but the challenge is that each entry of a column must be linked to a Wikipedia entry.
@@
Traditional statistical models need data assembled in a single consistent table, which is tackled by the data integration literature (Doan et al., 2012). Finding correspondences between columns across data sources is known as schema matching. Entity matching is a challenge common to data integration and natural language processing (NLP), where a string must be linked to an entity: a unique concept. For instance “Davinci” may denote the historical figure “Leonardo da Vinci”, but also OpenAI “Text-Davinci-003” GPT3 API. Entity matching must be robust to string variations, but most importantly it must account for the context in which an entity appears to disambiguate potential matches. In NLP, pretrained attention-based models, such as BERT (Devlin et al., 2019), have been crucial to capture the corresponding context.
These pretrained language models have actually been shown on tables to be useful to automate data normalization and integration tasks with few manually-supplied examples per table (Narayan et al., 2022). Deep learning, and more recently attention-based models, is progressing on tasks to structure databases, e.g. column typing, entity linking (Hulsebos et al., 2019; Deng et al., 2020).
We are interested in a different problem, capturing only implicit data structure and integration to enhance downstream machine learning task without any manual operation such as finding related sources. The problem is timely: researchers sharing this vision are assembling large corpora of tables (Hulsebos et al., 2023; Eggert et al., 2023). Yet, the small scale of most tables available and the variability between tabular datasets has so far made this vision elusive.
@@ 3.1 Graph Representation of Table Entities
The graph representation of table entities play a crucial role in facilitating the generalization power of CARTE. In general, a graph, $G$, consists of nodes and edges, where the former denote the entities and the latter denote the relations between the nodes. Graphs are useful for capturing intricate relations between the entities, and numerous graph-based models exploit such relations. The conventional usage of graphs for tables is to describe each row with a node and build edges based on the similarity of the features.
@@
On contrary, CARTE considers each of the row as a small graph, as shown in Figure 1. From a table with $k$ columns, CARTE represents each of the $i$-th instance as a graph, $G_{i}(X,E)$, where the components denote the node and edge features, respectively. The structure of $G_{i}(X,E)$ is a star-like graph with $k-p_{i}$ leaf-nodes, with $p_{i}$ as the number of columns with missing values for row $i$. On the resulting graphlet, each of the leaf-nodes are annotated by the cell values and their corresponding column names. To make these graphlets as viable inputs for neural networks, we initialize $X$ and $E$ by using a language model. For categorical values and column names, CARTE simply places a $d$-dimensional embedding that is generated from a language model. For numerical values, the features are initialized by the product of its value with the embedding of the corresponding column name. For instance, the node feature $X_{(239)}$ in Figure 1 is equal to $239\times E_{\text{(H index)}}$. Lastly, the center node is initialized with the mean of the leaflets, and will later serve as the readout-node that captures the overall information of the graph.
@@
Figure 1: Graphlet representation of tabular entities. From a table, CARTE represents each row as a star-like graphlet. Excluding for missing values, the leaf-nodes and the edges are annotated by the cell values and their corresponding column names. Then, CARTE initializes the features of nodes and edges with a language model. For numerical values the nodes are initialized by the elementwise product with its corresponding column features. For the center node, it is initially set as the average of the leaflets. It will later act as a readout that captures the overall information of the graphlet.
@@
There are several important aspects to be addressed. First, for tabular data, the context of an instance cannot be interpreted fully with the column names. In Figure 1, for example, it would be difficult to grasp the row (blue-box) only with the entries of ‘JMLR’, ‘15337928’, ‘$239$’. The column names ‘Title’, ‘ISSN’, and ‘H index’ clarifies that it is an instance of a journal. CARTE embodies the notion of context in tables through the nodes and edge, which can be exploited by its neural network architecture.
@@
Second, CARTE benefits from the use of language models on non-numerical entries, such as strings, categories, and names. With the language models in feature initialization, the graph transformation in CARTE does not require extensive intervention on discrete entries with data preprocessing or data-cleaning. Moreover, CARTE works with an open set of vocabularies. Problems of typos or wordings of the same meaning, such as ‘North America’ to ‘Northern America’, are readily resolved for CARTE.
@@
This naturally leads to our last point that the proposed graph representation is a powerful representation that can be generalized across heterogeneous tables. The graph transformation with CARTE synchronizes instances from different tables to the same graph domain, without requiring any schema matching for columns or entity matching for entries. Thus, learning process can be operated across many tables, which opens the door for pretraining or transfer.
@@ 3.2 Pretrained Model from a Large Knowledgebase
The pretraining process of CARTE can be summarized with Figure 2. From the large knowledge graph, we first extract small graphlets of entities that would work as inputs for CARTE. To enable the self-supervised scheme with a contrastive loss, we also generate truncated versions of selected graphlets and forge them as a batch. Through the training process, CARTE learns how to aggregate information based on the given context. In this subsection, we describe the specific details of the pretraining process.
Figure 2: CARTE pretraining process. From a large knowledge graph, CARTE begins by constructing graphlets and their positives of entities present in YAGO. The extracted samples are then fed into the CARTE neural network and trained with a self-supervised scheme. The resulting neural network learns how to aggregate information within the graphlets, which reflect how table entities can combine entries through the edges (columns)
@@
From the large knowledge graph of YAGO, we construct small graphlets of the entities that can be used as inputs for CARTE. To construct a suitable graphlet for an entity, we first extract its subgraph within a user-specified $k$-hop relations. To resemble the structure outlined in Figure 1, while benefiting from additional information through multiple hops, we set $k=2$ with a restriction in the maximum number of $1$-hop and $2$-hop relations as $100$ and $10$, respectively. However, unlike the graph structure shown in Figure 1 the center node would be the name of the entity (for instance ‘Louvre’). To account for the difference, we create an additional neighbor which is comprised of the name as its node and ‘has name’ as its relation. Finally, we initialize node and edge features with the procedure outlined in subsection 3.1 with the FastText embeddings (Mikolov et al., 2017) as the language model.
@@
To construct a batch sample of $N_{b}$, we first select which of the YOGO entities to be included in the batch as graphlets. For this, we employ a selection criteria where $0.9$ of $N_{b}$ are sampled from entities with $6$ or more $1$-hop relations and the remaining $0.1$ from the other subset. The main reason for such sampling scheme is that a large portion of entities in YAGO only have one or two $1$-hop relations, which seldomly come across for tabular data. Moreover, the value $6$ was selected so that the rough median of $1$-hop relations in the batch samples is $15$. To enable the self-supervised contrastive loss, we include positive samples, which are simply the truncation of original graphlets. Through the process, a random fraction (varying from $0.3$ to $0.7$) of the edges were deleted. An exemplary graphlet of ‘Louvre’ and its positive is shown in Figure 2.
@@
Figure 3 depicts the model structure of CARTE. At its basis, CARTE takes the classical Transformer encoder model in Vaswani et al. (2017), and adapts to a graph attentional network. The key component in CARTE’s architecture is the self-attention layer where both the node and edge features are considered in calculating the attention. The term ‘attention’ in the graph domain refers to the importance of neighbors from a specific node of interest (Velickovic et al., 2017). For graph representation of table entities, the attention translates to the importance of the entries for a given instance with the context supplemented by the column information.
@@
Next, we provide some mathematical details to see how CARTE leverages the the context for the self-attention mechanism. To bring some consistency in notations, we write vectors with an arrow on top $\vec{A}$, matrices in bold $\mathbf{A}$, and scalars $A$. For the ease of reading, we present the case of single-head attention layer, but keep in mind that it can easily be extended to the multi-head schemes of concatenating or averaging the attention outputs.
@@
For a graph with $N$ nodes, let $\vec{X}^{(l)}_{i}\in\mathbb{R}^{d}$ denote the feature of node $i$ and $\vec{E}^{(l)}_{ij}\in\mathbb{R}^{d}$ the feature of the edge directed from node $j$ to $i$. By design, graphlets for CARTE always holds the center node, which we fix its index at $1$. To obtain the representation from the attention layer, we first need to set query, key, and value, which are crucial elements for taking context into account. For query, it should be the vectors of what we are interested in, which are the nodes. Thus, we take the conventional approach and parameterize it with solely the node information. On the other hand, the key-value pairs should convey the elements that the neighboring nodes can offer. Therefore, we additionally parameterized them with the edge information. Keeping this in mind we set the three components as [^1] Here, we omit the superscripts indicating layers for $A_{ij}$, $e_{ij}$, and projection weights for $Q$, $K$, $V$ for clarity of presentation.:
$\vec{Q}_{i}$ $=\vec{X}^{(l)}_{i}\cdot\mathbf{W_{Q}}$
$\vec{K}_{ij}$ $=(\vec{X}^{(l)}_{i}\odot\vec{E}^{(l)}_{ij})\cdot\mathbf{W_{K}}$
$\vec{V}_{ij}$ $=(\vec{X}^{(l)}_{i}\odot\vec{E}^{(l)}_{ij})\cdot\mathbf{W_{V}}$
@@
where $\odot$ denotes the element-wise multiplication and $\mathbf{W_{Q}}$, $\mathbf{W_{K}}$, and $\mathbf{W_{V}}$ are trainable weights that reside on $\mathbb{R}^{d\times d}$. Here, the choice of element-wise product is motivated from a knowledge graph embedding technique in Balazevic et al. (2019); Cvetkov-Iliev et al. (2023). These works showed better performance with element-wise multiplication applied on edge and node features compared to translational operations. Following the scaled dot-product attention with the above three equations, the attention score of node $j$ from node $i$, $A_{ij}$, is derived as:
@@
where the calculation of $A_{ij}$ only takes the sum with respect to the connected neighbors of node $i$. This corresponds masking step, in which we take the inherent graph structure of the input. Notice that if we only considered the node features in $\vec{V}_{ij}$, then the attention score would only be the dot product (a scaled similarity) of projected word vectors of the entities. In such cases, the network would not be able to distinguish between the two cities, ‘Paris’ in France and ‘Paris’ in Texas. Through the attention layers, the outputs for nodes and edges become:
$\vec{X}_{i}^{(l+1)}$ $=\sigma_{X}(\sum_{j}A_{ij}\cdot\vec{V}_{ij})$ (2)
$\vec{E}_{ij}^{(l+1)}$ $=\sigma_{E}(E^{(l)}_{ij}\cdot\mathbf{W_{E}})$ (3)
where $\sigma$ denote the appropriate consecutive operations (see Figure 3). Finally, the consequent layers consist of the same attention layer, without the edge update, followed by the readout layer that extracts the representation of the center node ($i=1$). The outputs are then processed for the contrastive loss used in the pretraining. For more information on model specification and training details, please refer to Appendix A.1
@@
For the self-supervised contrastive loss, we adapted the framework outlined in Chen et al. (2020b). For CARTE, the original and one truncated graphlets are set as positives while the other graphlets in the batch are considered as negatives. The learning loss is then calculated based on the cosine similarity of the network outputs, accompanied by the InfoNCE loss (Oord et al., 2018).
Figure 3: The model architecture for CARTE The inputs of CARTE are graphs that contain node ($X$) and edge ($E$) features, which are both utilized self-attention layer (shown in grey). The attention layers focus on updating the node features based on the context embodied with the edge information. Moreover, the inherent graph structure of the input is reflected in the masking step. The consequent Aggregate & Readout layer consists of the same attention layer (without the edge update) followed by the extraction of the embedding on the center node. The outputs are then processed for the contrastive loss.

### v2

@@ Abstract
Pretrained deep-learning models are the go-to solution for images or text. However, for tabular data the standard is still to train tree-based models. Indeed, transfer learning on tables hits the challenge of data integration: finding correspondences, correspondences in the entries (entity matching) where different words may denote the same entity, correspondences across columns (schema matching), which may come in different orders, names… We propose a neural architecture that does not need such correspondences. As a result, we can pretrain it on background data that has not been matched. The architecture –CARTE for Context Aware Representation of Table Entries– uses a graph representation of tabular (or relational) data to process tables with different columns, string embedding of entries and columns names to model an open vocabulary, and a graph-attentional network to contextualize entries with column names and neighboring entries. An extensive benchmark shows that CARTE facilitates learning, outperforming a solid set of baselines including the best tree-based models. CARTE also enables joint learning across tables with unmatched columns, enhancing a small table with bigger ones. CARTE opens the door to large pretrained models for tabular data.
Machine Learning, ICML
@@ 1 Introduction
The wide availability of pre-trained models has greatly facilitated machine learning for various data modalities, for example with images (Simonyan & Zisserman, 2015) or texts (Devlin et al., 2019). These models can be downloaded from model hubs, embarking a lot of implicit information and transformations that unleashes the power of deep learning even on small datasets. This paradigm has led to the revolution of foundation models (Bommasani et al., 2021) such as large language models (Touvron et al., 2023). But this revolution has not happened for tabular data despite its huge importance for enterprise and institutional data. One roadblock is that integrating data from tables in the wild is often difficult, sometimes impossible. Different tables might not have any related data and when they do, data integration is a whole field of database research (Doan et al., 2012). One might need to solve correspondence across columns –schema matching– or across data sources with different naming conventions for entries –entity matching. For lack of matching schemas and entities, pretraining across tables in the wild has not been possible. Without pretraining, deep learning is less practical, and tree-based methods are often preferable (Grinsztajn et al., 2022).
Here we introduce a learning architecture that learns across tables without schema and string matching. The key is to represent tables with a graph and all symbols with embeddings (for column names and table entries). The architecture, dubbed CARTE (Context-Aware Representation of Table Entries), is pretrained on a large knowledge base, to capture information on a vast amount of entities and relations. It can then be fine-tuned on a given downstream task, helping learning even in few shot settings. It can also be used for joint learning across multiple tables, enriching a target table with weakly related sources. CARTE brings a sizable performance gain, outperforming markedly a set of 42 solid baselines (including the best tree-based methods and various feature engineering). It benefits particularly tables with string entries, frequent in applications but seldom present in machine learning benchmarks.
Section 2 presents related work; section 3 describes the CARTE architecture and training procedures; and section 4 provides an extensive empirical study across many tabular datasets, benchmarking the settings of a single downstream table as well as multiple related ones.
@@ 2 Related Works
Tables are central to many applications. As a result, numerous deep learning methods tailored for this modality have been proposed (Abutbul et al., 2020; Arik & Pfister, 2020; Popov et al., 2019; Gorishniy et al., 2023b; Somepalli et al., 2021). However, they typically under-perform tree-based methods (Grinsztajn et al., 2022; Shwartz-Ziv & Armon, 2021; Gardner et al., 2022). While McElfresh et al. (2023) argue that neural networks perform well on certain types of tables, and promising architectures are continuously published (Gorishniy et al., 2023a; Chen et al., 2023), the difficulty of improving over tree-based methods suggests that deep learning must bring something more to the fight, such as background knowledge.
@@
Transfer learning mostly focuses on the “conventional” settings of transferring across datasets with the same features, i.e. columns. Somepalli et al. (2021) demonstrates pre-training on a larger unlabeled version of the table, while Levin et al. (2023) argues that transfer learning bridges the gap between deep learning and tree-based models when there are few data points but large related datasets, as in their medical setting. They consider new or missing features in the downstream table, but require exact matching between most features.
@@
TabPFN (Hollmann et al., 2023) made headway in pre-training models for tabular learning: it uses a transformer model pre-trained on large amounts of synthetic data to capture the inductive biases of tabular data, leading to strong performance on small datasets, though it has no dedicated handling for categorical columns, a challenge of tables where trees historically shine. Large language models (LLMs) can also work as pretrained models for tabular data. In TabLLM (Hegselmann et al., 2023), tabular data are represented as a set of tokens which are leveraged to fine-tune an LLM. However, the difficulty of handling numerical values in LLMs makes them a suboptimal choice compared to trees or TabPFN.
@@
One challenge of tables –much more tackled by the database literature than the machine learning one– is that many of the entries are discrete, represented as strings. Cerda & Varoquaux (2022) created string-based representations that facilitate learning. The TableVectorizer in Skrub (2024) uses these heuristically to turn columns of different types into numerical matrices well suited for learning. KEN (Cvetkov-Iliev et al., 2023) is another approach to embed table entities closer to our goals of pretraining. It provides embeddings of all entities in a knowledge graph, capturing the information in a source such as Wikipedia. These embeddings facilitate learning, but the challenge is that each entry of a column must be linked to a Wikipedia entry, an entity matching task.
@@
Traditional statistical models need data assembled in a single consistent table, a task tackled by the data integration literature (Doan et al., 2012). Finding correspondences between columns across data sources is known as schema matching. Entity matching is a challenge common to data integration and natural language processing (NLP), where a string must be linked to an entity: a unique concept. For instance “Davinci” may denote the historical figure “Leonardo da Vinci”, but also OpenAI “Text-Davinci-003” GPT3 API. Entity matching must be robust to string variations, but most importantly it must account for the context in which an entity appears to disambiguate potential matches. In NLP, pretrained attention-based models, such as BERT (Devlin et al., 2019), have been crucial to capture the corresponding context.
These pretrained language models also are useful on tables to automate data normalization and integration tasks with few manually-supplied examples (Narayan et al., 2022). Deep learning, and recently attention-based models, is progressing on tasks to structure databases, e.g., column typing, entity linking (Hulsebos et al., 2019; Deng et al., 2020).
We are interested in a different problem: rather than explicit matching, at the entity or schema level, we aim to capture only implicit data structure and integration to enhance downstream machine learning task without any manual operation such as finding related sources. The problem is timely: researchers sharing this vision are assembling large corpora of tables (Hulsebos et al., 2023; Eggert et al., 2023). Yet, the small scale of most tables available and the variability between tabular datasets has so far made this vision elusive.
@@ 3.1 Graph Representation of Table Entities
The graph representation is crucial for facilitating the generalization of table entities. In general, a graph, $G$, consists of nodes and edges, where the former denote the entities and the latter denote the relations between the nodes. Graphs are useful for capturing relational information between entities, and graph deep learning is promising for relational databases (Fey et al., 2023). CARTE considers each of the row as a small graph, as shown in Figure 1. From a table with $k$ columns, CARTE represents each of the $i$-th instance as a graph, $G_{i}(X,E)$, where the components $X$ and $E$ denote the node and edge features, respectively, embeddings in $\mathbb{R}^{d}$. The structure of $G_{i}(X,E)$ is a star-like graph with $k-p_{i}$ leaf-nodes, with $p_{i}$ as the number of columns with missing values for row $i$. On the resulting graphlet, each of the leaf-nodes are annotated by the cell values and their corresponding column names. To make these graphlets viable inputs for neural networks, we initialize $X$ and $E$ by using a language model. For categorical values and column names, CARTE simply places a $d$-dimensional embedding that is generated from a language model. For numerical values, the features are initialized by the product of its value with the embedding of the corresponding column name. For instance, the node feature $X_{(239)}$ in Figure 1 is equal to $239\times E_{\text{(H index)}}$. Lastly, the center node is initialized with the mean of the leaflets, and will later serve as the readout node that captures the overall information of the graph.
@@
Figure 1: Graphlet representation of tabular entities. From a table, CARTE represents each row as a star-like graph. Excluding for missing values, the leaf-nodes and the edges are annotated by the cell values and their corresponding column names. Then, CARTE initializes the features of each with a language model. The nodes of numerical values are initialized by the elementwise product with its corresponding column feature. For the center node, it is initially set as the average of the leaflets. It will later act as a readout that captures the overall information of the graphlet.
@@
This design serves several purposes. First it represents context: for tabular data, an entry is best interpreted accounting for its column’s name. In Figure 1, for example, it would be difficult to grasp the row (blue-box) only with the entries of ‘JMLR’, ‘15337928’, ‘$239$’. The column names ‘Title’, ‘ISSN’, and ‘H index’ clarifies that it is an instance of a journal. CARTE represents context in tables through the nodes and edge, exposed to its neural network architecture. This representation also bridges tables with different column order, or more generally different columns.
@@
Second, CARTE uses language models on non-numerical entries, such as strings, categories, and names. Thus, the graph transformation in CARTE does not require any intervention on discrete entries, as opposed to typical data preprocessing or cleaning (deduplication, categorical encodings) used on strings. Moreover, CARTE works with an open set of vocabularies. Problems of typos or wordings of the same meaning, such as ‘North America’ to ‘Northern America’, are readily resolved for CARTE.
@@
Together, these features of the proposed graph representation enable generalization across heterogeneous tables. CARTE’s graph transformation puts in the same graph domain instances from different tables, without requiring any schema matching for columns or entity matching for entries. Thus, learning process can operate across many tables, which opens the door for pretraining or transfer.
@@ 3.2 Pretrained Model from a Large Knowledge Base
In this subsection, we describe the pretraining process of CARTE, summarized in Figure 2. From the knowledge graph, we first extract small graphlets of entities suitable as inputs for CARTE. For self-supervised learning with a contrastive loss, we add to the batch truncated versions of selected graphlets. Through the training process, CARTE learns to aggregate information based on the given context.
Figure 2: CARTE pretraining process. From a large knowledge graph, CARTE begins by constructing graphlets and their positives variants. The extracted samples are then fed into the CARTE neural network and trained with a self-supervised scheme. The neural network learns to aggregate information within the graphlets, which reflect the combination of table entries across columns (edges).
@@
From the large knowledge graph of YAGO, we construct small graphlets of the entities that can be used as inputs for CARTE. To construct a suitable graphlet for an entity, we first extract its subgraph within a user-specified $k$-hop relations. To resemble the structure outlined in Figure 1 while benefiting from additional information through multiple hops, we set $k=2$, but restricting the maximum number of $1$-hop and $2$-hop relations to $100$ and $10$ respectively. Graphlets from tables (Figure 1) have as center node a token for the row, while the knowledge-graph procedure could use the entity name (for instance ‘Louvre’). To avoid a difference, we use a token as a center node with an additional neighbor which is comprised of the name as its node and ‘has name’ as its relation. Finally, as in subsection 3.1, we initialize node and edge features using FastText embeddings (Mikolov et al., 2017) as the language model.
@@
To construct a batch sample of size $N_{b}$, we first select which of the YAGO entities to include, generating the corresponding graphlets. For this, we sample $0.9$ of $N_{b}$ from entities with $6$ or more $1$-hop relations and the remaining $0.1$ from the other subset. The main reason for such sampling scheme is that a large portion of entities in YAGO only have one or two $1$-hop relations, while tabular data typical has more (more columns). Moreover, the value $6$ was selected so that the rough median of $1$-hop relations in the batch samples is $15$. To enable the self-supervised contrastive loss, we include positive samples, which are simply truncations of original graphlets: deleting a random fraction (varying from $0.3$ to $0.7$) of the edges. Figure 2 gives an exemple graphlet of ‘Louvre’ and its positive.
@@
Figure 3 depicts the model structure of CARTE. At its basis, CARTE takes the classical Transformer encoder model of Vaswani et al. (2017), and adapts to a graph attentional network. A key component in CARTE’s architecture is a self-attention layer computing attention from both node and edge features. In graph models, attention modulates the importance of neighbors for a given node of interest (Velickovic et al., 2017). For table entries, it translates to the importance of the entries for a given instance with the context supplemented by the column information.
Figure 3: CARTE architecture The inputs of CARTE are graphs that contain node ($X$) and edge ($E$) features, both used in self-attention layers (shown in grey). The attention layers update node features using the context embodied with the edge information; the graph structure of the input is reflected by attention masks. The Aggregate & Readout layer consists of the attention layer (without the edge update) followed by feature extraction on the center node. The outputs are then processed for the contrastive loss.
@@
We now detail on CARTE’s attention mechanism used to capture context and relations. For consistent notations, we write vectors with an arrow on top $\vec{A}$, matrices in bold $\mathbf{A}$, and scalars $A$. To ease reading, we present a single-head attention layer, but it can easily be extended to multi-head schemes of concatenating or averaging the attention outputs.
@@
For a graph with $N$ nodes, let $\vec{X}^{(l)}_{i}\in\mathbb{R}^{d}$ denote the feature of node $i$ and $\vec{E}^{(l)}_{ij}\in\mathbb{R}^{d}$ the feature of the edge directed from node $j$ to $i$. By design, graphlets for CARTE always hold the center node, which we denote with an index $i=1$. The representation from the attention layer is a function of query, key, and value, crucial elements to account for context. The query is the vectors corresponding to our values of interest: the nodes. Thus, we take the conventional approach and parameterize it with solely the node information. On the other hand, the key-value pairs should convey the elements that the neighboring nodes can offer. Therefore, we add edge information in the corresponding parameterization. With this in mind we set the three components as [^1] Here, we omit the superscripts indicating layers for $A_{ij}$, $e_{ij}$, and projection weights for $Q$, $K$, $V$ for clarity of presentation.:
Query: $\vec{Q}_{i}$ $=\vec{X}^{(l)}_{i}\cdot\mathbf{W_{Q}}$ (1)
Key: $\vec{K}_{ij}$ $=(\vec{X}^{(l)}_{i}\odot\vec{E}^{(l)}_{ij})\cdot\mathbf{W_{K}}$ (2)
Value: $\vec{V}_{ij}$ $=(\vec{X}^{(l)}_{i}\odot\vec{E}^{(l)}_{ij})\cdot\mathbf{W_{V}}$ (3)
@@
where $\odot$ denotes the element-wise multiplication and $\mathbf{W_{Q}}$, $\mathbf{W_{K}}$, and $\mathbf{W_{V}}$ are trainable weights that reside on $\mathbb{R}^{d\times d}$. Here, the choice of element-wise product is motivated from a knowledge graph embedding technique in Balazevic et al. (2019); Cvetkov-Iliev et al. (2023). These works showed that modeling relations (i.e. column name) as element-wise multiplication on the node vectors works best, compared to e.g., vector additions. Following the scaled dot-product attention with the above three equations, the attention score of node $j$ from node $i$, $A_{ij}$, is derived as:
@@
where the calculation of $A_{ij}$ only takes the sum with respect to the connected neighbors of node $i$. This corresponds to a masking step, which takes the graph structure of the input. Accounting for the relation type (i.e. column name) in the attention scores ($\vec{E}$ in eq 2 and 3) is important for proper re-contextualization of the entries, that is to capture their meaning. For instance, an entry “George Bush” may denote the 41st or 43rd US presidents, an aircraft carrier… The ambiguity is raised by the relation (“George Bush”, “son of”, “George Bush”), however capturing fully the information does require knowing the nature of the relation, as “father of” would lead to a different entity resolution. Ablations reveal the importance of attention (Appendix C.3).
Outputs of the attention layers are, for nodes and edges:
transformed entry $\vec{X}_{i}^{(l+1)}$ $=\sigma_{X}\bigl{(}\sum_{j}A_{ij}\cdot\vec{V}_{ij}\bigr{)}$
transformed relation $\vec{E}_{ij}^{(l+1)}$ $=\sigma_{E}(E^{(l)}_{ij}\cdot\mathbf{W_{E}})$
where $\sigma$ denote the appropriate consecutive operations (see Figure 3). The final layers consist of the attention layer without the edge update, followed by the readout layer that extracts the representation of the center node. For pretraining, the outputs are then processed for the contrastive loss. Appendix A.1 details model specification and training.
@@
For the self-supervised contrastive loss, we adapt the framework of Chen et al. (2020b). The original graphlet and one truncation are set as positives while the other graphlets in the batch are considered as negatives. The learning loss is then based on the cosine similarity of the network outputs, fed in the InfoNCE loss (Oord et al., 2018).